Define p-adic number with example.


1
Expert's answer
2021-06-01T13:09:24-0400

Every natural number m defines a sequence:

"a_n\equiv m(mod\ p^n)" , p is a prime number

and can therefore be regarded as a p-adic integer.

For example, 35 as a 2-adic integer would be written as the sequence

"(1,3,3,3,3,35,35,35,...)"

"a_1=35(mod\ 2)=1"

"a_2=35(mod\ 2^2)=3"

"a_2=35(mod\ 2^3)=3"

etc.